# Re-Examine Distantly Supervised NER: A New Benchmark and a Simple Approach

### v1

@@ 4.3 Curriculum-based PU Learning
Specifically, we adopt the Conf-MPU loss function, proposed by Zhou et al. (2022), as the backbone PU loss function in the curriculum-based training. Conf-MPU loss function has been shown to be more robust to PU assumption violation in practice. Instead of having entity confidence score $\lambda(x)$ estimated by another binary PU model, the only difference we make is to reuse the voters trained in Section 4.1 to ensemble the confidence score for each token $x$. We use the soft-label ensemble as
@@
where $\boldsymbol{\Pr}(x)$ is the ensemble probability distribution over all classes.
The confidence score of a token $x$ being an entity token is then calculated as
@@
For the neural network of the NER classifier, we choose the same structure with voters, which is defined at the beginning of Section 4.
@@ 4.4 Self-Training
Several studies Liang et al. (2020); Peng et al. (2019); Meng et al. (2021) have shown that self-training can effectively upgrade the performance of a trained DS-NER model. We apply the self-training method in Meng et al. (2021), which uses soft labels to conduct self-training and uses a masked language model to conduct contextual data augmentation simultaneously. Self-training is used directly after CuPUL, and we call the final classifier after self-training CuPUL+ST.
5 Experimental Study
@@ 5.1 Baseline Methods
We use two groups of baseline methods.
Fully supervised methods. We include a fully supervised NER method based on the RoBERTa-base model Liu et al. (2019) as an upper-bound performance reference.
Distantly-supervised methods. First, we report distant supervision results as KB-Matching. We classify DS-NER methods into three groups. 1) DS-NER without Self-training consists of AutoNER Shang et al. (2018), Conf-MPU Zhou et al. (2022), RoBERTa-ES Liang et al. (2020). CuPUL is directly comparable with these methods. 2) DS-NER with Self-training includes BOND Liang et al. (2020), RoSTER Meng et al. (2021), SCDL Zhang et al. (2021b) and ATSEN Qu et al. (2023). These methods apply teach-student or training augmentation steps to further boost the DS-NER performance. 3)Span-based DS-NER models includes Top-Neg Xu et al. (2023), CLIM Li et al. (2023) and SANTA Si et al. (2023), and all other methods use sequence-based models. Previous work Li et al. (2023) shows that span-based NER models often outperform sequence-based NER methods in terms of effectiveness, albeit at the cost of increased algorithmic complexity. More details of baselines can be found in Appendix C.
5.2 Experiment Settings
We use the pre-trained RoBERTa as the backbone model for both the Voter and NER classifier [^3] We will release code upon paper acceptance.. For open-domain datasets, we use roberta-base [^4] https://huggingface.co/roberta-base. For bio-domain datasets, we use biomed-reberta-base [^5] https://huggingface.co/allenai/biomed_roberta_base. We report single-run results for the model performance and the random seed is set to 42. We employ PyTorch [^6] https://pytorch.org/ and conduct all experiments on a server with a Tesla A100 GPU (32G).
@@ 5.3 QTL Experiments
We discuss the performance of CuPUL on QTL dataset in this section.
5.3.1 Re-Examine Baseline Methods
In our exploration of baseline methods, we encountered several practical problems. For example, BOND [^7] https://github.com/cliang1453/BOND uses the test set for hyperparameter tuning during training, so we made modifications to their published code to tune hyperparameters on the validation set. RoSTER lacks tuning guidelines in the paper, so we used the reported hyperparameters from Meng et al. (2021) directly and only altered the max_seq_length to 300, since the sentence length of QTL is much longer. Finally, CLIM, Top-Neg, and SANTA lack runnable published code, so we did not include them in the evaluation. We also encountered some other practical issues, which are discussed in the Appendix E.
5.3.2 Results on QTL
@@
CuPUL+ST 61.83 (58.65/65.38) 75.82 (72.96/78.91)
Table 1: Performance on QTL dataset: F1 Score (Precision/Recall) (in %). The best results are in bold, second best results are in underline.
Table 1 presents strict and relaxed precision, recall, and F1 scores for all methods on the QTL dataset. Note that CuPUL without curriculum learning (CuPUL-curr) is essentially equivalent to Conf-MPU when there is one entity type. KB matching reveals that QTL annotations suffer from low recall but have relatively high precision, creating a notable imbalance between these metrics. We observe that DS-NER baselines without self-training have limited recall improvement, resulting in weak performance. DS-NER baselines with self-training improve recall, but CuPUL and CuPUL+ST can further boost the recall, significantly outperforming all baseline methods. Specifically, strict F1 and relaxed F1 of CuPUL+ST outperform the runner-up by 8.75% and 10.25%, respectively.
@@ 5.4 Benchmark Experiments
We discuss the performance of CuPUL on existing benchmark datasets in this section.
@@ 5.4.1 Datasets and Metrics
Datasets: We conduct experiments on six existing benchmark datasets including CoNLL03 Liang et al. (2020), Twitter Liang et al. (2020), OntoNotes5.0 Liang et al. (2020), Wikigold Liang et al. (2020), Webpage Liang et al. (2020), and BC5CDR Shang et al. (2018). The first five are open-domain datasets, and BC5CDR is the bio-medical domain. More details and the statistics of these datasets are summarized in Appendix B.
Settings: For the benchmark dataset, we use small subsets of the validation set to tune the hyperparameters including learning rate, epochs, etc, to simulate the real-life DS-NER application scenarios. Detailed settings and statistics of the validation set can be found in Appendix D.
5.4.2 Results on Benchmark Datasets
Method CoNLL03 Twitter OntoNotes5.0 Wikigold Webpage BC5CDR
Fully Supervised
${}^{\#}$ 90.11 (89.14/91.10) 52.19 (51.76/52.63) 86.20 (84.59/87.88) 86.43 (85.33/87.66) 72.39 (66.29/79.73) 90.99 (${}^{\dagger}$
Span-based DS-NER models
${}^{\Diamond}$ 86.59 (86.25/86.95) - 69.72 (69.24/70.21) - 71.79 (78.40/66.72) 79.23 (81.74/76.88)
${}^{\Diamond}$ 80.55 (81.07/80.23) 52.86 (52.30/53.55) - - - 80.39 (82.09/78.90)
${}^{\Diamond}$ 85.4 (-/-) 53.8 (-/-) 69.6 (-/-) 70 (-/-) 67.9 (-/-) -
DS-NER without Self-training
${}^{\#}$ 71.40 (81.13/63.75) 35.83 (40.34/32.22) 59.51 (63.86/55.71) 47.76 (47.90/47.63) 52.45 (62.59/45.14) 64.32 (${}^{\dagger}$
${}^{\#}$ 67.00 (75.21/60.40) 26.10 (43.26/18.69) 67.18 (64.63/69.95) 47.54 (43.54/52.35) 51.39 (48.82/54.23) 79.99 (${}^{\dagger}$
${}^{\#}$ 75.61 (83.76/68.90) 46.61 (53.11/41.52) 68.11 (66.71/69.56) 51.55 (49.17/54.50) 59.11 (60.14/58.11) 73.66 (${}^{\dagger}$
${}^{\dagger}$ 79.16 (78.58/79.75) - - - - 77.22 (${}^{\dagger}$
CuPUL 85.09 (84.64/85.53) 54.34 (54.47/54.20) 68.06 (66.31/69.91) 70.53 (67.06/74.39) 73.10 (74.65/71.62) 81.57 (77.02/86.70)
DS-NER with Self-training
${}^{\#}$ 81.15 (82.00/80.92) 48.01 (53.16/43.76) 68.35 (67.14/69.61) 60.07 (53.44/68.58) 65.74 (67.37/64.19) -
[…]
${}^{\ddagger}$ 83.69 (87.96/79.82) 51.10 (59.87/44.57) 68.61 (67.49/69.77) 64.13 (62.25/66.12) 68.47 (68.71/68.24) -
${}^{\ddagger}$ 85.59 (86.14/85.05) 52.46 (62.32/45.30) 68.95 (66.97/71.05) - 70.55 (71.08/70.55) -
Table 2 presents the overall span-level precision, recall, and F1 scores for all methods on benchmark datasets. Note that RoSTER was tested on a different version of the OntoNotes5.0 dataset Meng et al. (2021). Therefore, we exclude its reported results in Table 2 for a fair comparison. Note that the results of baseline methods are obtained from previous papers, and we did not rerun the experiments. The baseline methods used different validation sets for parameter tuning, which are much larger. We have the following observations.
Span-based DS-NER models. This group of baselines use span-based models. It is unfair to compare them with sequence-based models directly, so we just put it in Table 2 to inspire potential insights.
DS-NER Without Self-training. From Table 2, it is obvious that KB-Matching generally exhibits low recall and on four of the benchmark datasets, low precision as well. In contrast, noise-aware DS-NER models, like CuPUL, significantly outperform KB-Matching. This is confirmed in the table where CuPUL achieves the best F1 scores on four out of six datasets compared to all DS-NER models and displays competitive results on the OntoNotes5.0 dataset. On the OntoNotes5.0 dataset, almost all noise-aware DS-NER models have similar performances, implying that distant annotations may contain biases difficult for the models to address.
DS-NER With Self-training. The results for CuPUL+ST shown in Table 2 further verify that adding a self-training phase tends to enhance overall performance. When compared with DS-NER models that incorporate self-training, CuPUL+ST demonstrates superior performance on five datasets. This indicates that the CuPUL model benefits from the self-training approach, making it a versatile and effective tool for various datasets. Surprisingly, CuPUL and CuPUL+ST achieve higher F1 scores than the supervised baseline (RoBERTa) on the Twitter and Webpage datasets. This could be due to inconsistent distributions of training and test data for these datasets, and the supervised learning may have overfitted the training set due to its limited size.
5.5 Further Analysis
To further validate the effectiveness of CuPUL, we conducted additional analyses, including an ablation study. We are unable to use the QTL dataset for this purpose because the train set for QTL lacks ground truth, preventing us from analyzing the curriculum quality during training process. Specifically, we utilized open-domain datasets such as Wikigold and Twitter as representatives.
5.5.1 Difficulty Score Estimation
Figure 2: Distant Label Quality Token Level Positive Error Rate and Mean Difficulty Scores for Each Curriculum on Wikigold Dataset.
Figure 3: Distribution of the Difficulty Scores for Labeled Positives on Wikigold Dataset
For CuPUL, one assumption adopted is that difficulty scores can reflect the quality of distant supervision, where “easier” tokens have “cleaner” labels. To validate this assumption and evaluate the quality of the difficulty score estimation, we examine the correlation between the difficulty scores and the quality of distant labels. We use Wikigold as the testbed, and the results are illustrated in Figure 2.
For each training curriculum, we compute the token-level positive error rate (positive errors include false positives and positive type errors), and plot the rate for each curriculum using the left y-axis in Figure 2. We also compute the average difficulty scores for tokens in each curriculum shown with the right y-axis in Figure 2.
We can see that as the number of curricula increases, the average token difficulty scores and positive error rate have a clear increase. This illustrates a strong negative correlation between the difficulty scores and the quality of distant labels. Specifically, as the difficulty score increases, the quality of the distant labels significantly decreases. This result validates our assumption that “easy” data have cleaner labels and “hard” data have noisier labels. The clean data can initialize the model from a better starting point and improve the model’s robustness to noise in the latter curricula.
Another important assumption we adopt for the design of curricula is that the difficulty scores follow a long-tail distribution. We illustrate the distribution of difficulty scores estimated on the Wikigold dataset in Figure 3. It clearly demonstrates the long-tail phenomenon, with most tokens having low difficulty scores. This phenomenon can be observed in other datasets. Due to the space limit, we omit the plots for other datasets.
5.5.2 Ablation Study
Curriculum Learning To evaluate the effectiveness of curriculum learning in CuPUL, we compare it with two variations of itself. First, we use the five voters trained using positive and sampled negative examples and take the average of their soft label predictions as the result. The results are shown as voter ensemble in Table 3. Second, we include the result of CuPUL-curr from Table 2 since it is another variation. To evaluate the effectiveness of the Conf-MPU loss estimation for curriculum learning in CuPUL, we use the regular loss estimation, which considers unlabeled tokens as non-entity tokens, denoted as w/o Conf-MPU in Table 3.
Our analysis reveals the critical role of each component, as removing any of them results in a significant drop in the F1 score. Compared CuPUL-curr with w/o Conf-MPU, we find that CuPUL-curr consistently achieves higher recall. This is attributed to Conf-MPU primarily addressing false positives and partial false positives Zhou et al. (2022), leading to more tokens being predicted as entities, thereby enhancing recall. Conversely, w/o Conf-MPU exhibits higher precision since it tackles both false positives and positive type errors. Addressing positive type errors benefits both precision and recall, but the increase in precision is more pronounced compared to CuPUL-curr.
Method Wikigold Twitter
Precision Recall F1 Precision Recall F1
CuPUL 67.06 74.39 70.53 54.47 54.20 54.34
w/o Curriculum Learning
voter ensemble 56.88 74.88 64.65 35.52 49.52 41.37
CuPUL-curr 58.89 76.18 66.43 47.48 53.07 50.12
w/o Conf-MPU 59.31 75.86 66.57 58.91 47.04 52.53
6 Conclusion and Future Work
In this paper, we introduce a real-life DS-NER application in the animal science domain, named QTL, and we reveal the limitations of current DS-NER methods in practical DS-NER settings on QTL. To solve this issue, we propose a simple yet effective token-level curriculum-based PU learning (CuPUL) method, which strategically sequences the training data from “easy” to “hard”. Our experiments show that CuPUL not only mitigates the adverse effects of noisy labels but also achieves state-of-the-art DS-NER on many datasets. Through CuPUL, we demonstrate the effectiveness of curriculum learning in improving the performance of DS-NER systems.
Limitations
Ethics Statement
We comply with the ACL Code of Ethics.
Appendix
.1 Curriculum Learning
Curriculum learning was first proposed by Bengio et al. (2009) under the assumption that learning with reordering from “easy” samples to “hard” samples would boost performance. It has been applied in various applications, including neural machine translation Zhou et al. (2020); Platanios et al. (2019); Zhou et al. (2020); Wang et al. (2018), relation extraction Huang and Du (2019), reading comprehension Tay et al. (2019), natural language understanding Xu et al. (2020) and named entity recognition Jafarpour et al. (2021); Lobov et al. (2022); Wenjing et al. (2021).
Appendix A Discussion of Loss Function
[…]
Another commonly used loss function is mean absolute error (MAE):
Figure 5: Screenshot for online annotation tool teamtat.
Appendix B Datasets
To annotate the QTL dataset, domain experts use an online tool named teamtat [^8] https://www.teamtat.org/. The screenshot of the tool shows in Figure 5.
  • BC5CDR comes from the biomedical domain. It consists of 1,500 articles, containing 15,935 Chemical and 12,852 Disease mentions.
The statistics of the baseline datasets are shown in Table 4.
Appendix C Baselines
RoBERTa-ES Liang et al. (2020) trains a NER model using a RoBERTa-base model and adopts early stopping to prevent the model from overfitting to noisy distant labels.
Appendix D Experiment Settings
For datasets CoNLL03, OntoNotes5.0, Webpage, Twitter, Wikigold, QTL and BC5CDR, the maximum sequence length is set as 150, 230, 120, 160, 120, 180, 280 respectively, to ensure the algorithm works correctly. Other parameters shows in Table 5. OntoNote5.0 and Twitter have more entity types, so we choose relatively small $\gamma$ values. We apply cross-entropy loss to OntoNotes5.0 and Twitter since they have more entity types and apply MAE loss to CoNLL03, Webpage, and Wikigold. For all the datasets, we train them with a batch size of 32 sentences and apply Adam optimizer Kingma and Ba (2014). For all the datasets, the number of voters $K$ and the number of curricula $C$ are set as 5 and 5, respectively. The curriculum selective factor $\tau$ is set to 0.5 for all the datasets. We use the same random seeds for all datasets.
To follow the real-life setting of DS-NER, we use the small subset of validation to adjust the learning rate on benchmark datasets. Other hyper-parameters are set according to data statistics (show in Table 4). For QTL dataset, baselines are reproduced using their published codes. Baselines CLIM, SANTA, and Top-Neg are excluded due to reproduction obstacles. Hyperparameters of baseline methods and CuPUL are tuned on the validation set.

### v3

@@ 4.3 Curriculum-based PU Learning
Specifically, we adopt the Conf-MPU loss function, proposed by Zhou et al. (2022), as the backbone PU loss function in the curriculum-based training.
Details of Conf-MPU can be found in Appendix B. Instead of having entity confidence score $\lambda(x)$ estimated by another binary PU model, the only difference we make is to reuse the voters trained in Section 4.1 to ensemble the confidence score for each token $x$. We use the soft-label ensemble as
@@
where $\boldsymbol{\Pr}(x)$ is the ensemble probability distribution over all classes. The confidence score of a token $x$ being an entity token is then calculated as
@@
For the neural network of the NER classifier, we choose the structure described at the beginning of Section 4.
@@ 4.4 Self-Training
Several studies Liang et al. (2020); Peng et al. (2019); Meng et al. (2021) have shown that self-training can effectively upgrade the performance of a trained DS-NER model. We apply the self-training method in Meng et al. (2021), which uses soft labels to conduct self-training and a masked language model to conduct contextual data augmentation simultaneously. Self-training is used directly after CuPUL, and we call the classifier with self-training “CuPUL+ST”.
5 Experimental Studies
@@ 5.1 Baseline Methods
We use feasible methods mentioned in Section 3 as baseline methods. First, we report distant supervision results as KB-Matching. We classify feasible DS-NER methods into two groups. 1) DS-NER without Self-training consists of AutoNER Shang et al. (2018) and Conf-MPU Zhou et al. (2022). CuPUL is directly comparable with these methods. We also include an ablation version of CuPUL (CuPUL-curr), which removes Curriculum Learning, as a baseline. 2) DS-NER with Self-training includes BOND Liang et al. (2020), RoSTER Meng et al. (2021), SCDL Zhang et al. (2021b) and ATSEN Qu et al. (2023) and DesERT Wang et al. (2023a). These methods apply teach-student or training augmentation steps to further boost the DS-NER performance. CuPUL+ST is directly comparable with these methods.
To ensure a fair comparison, we made some necessary code modifications to the baseline methods. For Conf-MPU, we updated the encoding model to RoBERTa. For BOND, SCDL, ATSEN, and DesSERT, we modified the hyperparameter tuning process to use the validation set instead of the test set. Early stopping is used to select the inference model. RoSTER uses fixed parameters, but the max_seq_length did not meet the requirements for some datasets, so we adjusted it accordingly. Specific parameters details are in Appendix F.
@@ 5.2 QTL Experiments
Evaluation Metrics: Due to the annotation limitation and the fact that none of the DS-NER methods can handle discontinued spans, we include relaxed Precision, Recall, and F1 scores to evaluate the performance on the QTL dataset, in addition to the strict span-level Precision, Recall, and F1 scores used in previous studies. For relaxed metrics, it deems a predicted span correct if there is at least one overlapping word with the ground truth annotation. According to the curator’s feedback, the relaxed metrics can meet the practical need as identifying potential entities is more important than identifying precise boundaries.
@@
DesERT 54.41 (69.20/44.83) 64.23 (82.41/51.50)
CuPUL+ST 58.87 (58.28/59.47) 73.57 (73.07/74.08)
Table 4: Performance on QTL dataset: F1 Score (Precision/Recall) (in %). The best results are in bold.
Method CoNLL03 Twitter OntoNotes5.0 Wikigold Webpage BC5CDR
DS-NER Without Self-training
KB-Matching * 71.40 35.83 59.51 47.76 52.45 64.32
AutoNER * 67.00 26.10 67.18 47.54 51.39 79.99
Conf-MPU $\dagger$ 82.39 43.21 66.04 66.58 63.32 80.06
CuPUL-curr 83.18 50.12 67.76 66.43 65.15 79.29
CuPUL 85.09 54.34 68.06 70.53 73.10 80.19
DS-NER With Self-training
RoSTER   85.40   43.91   69.10   58.34   56.80   79.78
BOND $\dagger$ 79.89 45.98 66.86 57.81 48.76 76.91
* 81.15 48.01 68.35 60.07 65.74 -
SCDL $\dagger$ 82.47 44.76 68.50 47.62 41.29 77.72
* 83.69 51.10 68.61 64.13 68.47 -
ATSEN $\dagger$ 79.39 49.38 68.22 60.72 43.03 79.95
* 85.59 52.46 68.95 - 70.55 -
DesERT $\dagger$ 80.57 48.21 67.94 60.32 62.88 78.21
* 86.95 52.26 69.17 65.99 72.73 -
CuPUL+ST 86.64 54.78 68.20 70.19 74.48 80.87
Table 5: Performance on benchmark datasets with small validation sets: F1 Score (in %). * marks the results reported from the original papers, and $\dagger$ marks the results we run. The best results are in bold. Data in gray font are NOT used for comparative analysis as they were tuned using either the test set or an large validation set. We include these only to contrast our re-run results with previous works.
Table 4 presents the results for all methods on the QTL dataset. Note that CuPUL without curriculum learning (CuPUL-curr) is essentially equivalent to Conf-MPU when there is one entity type. KB matching reveals that QTL annotations suffer from low recall but have relatively high precision. We observe that DS-NER baselines without self-training have limited recall improvement, resulting in weak performance. DS-NER baselines with self-training improve recall compared to AutoNER, but still generally under-perform compared to PU-based methods. CuPUL+ST can further boost the recall compared to CuPUL, significantly outperforming all baseline methods. Specifically, strict F1 and relaxed F1 of CuPUL+ST outperform the runner-up by 5.79% and 8.00%, respectively.
@@ 5.3 Benchmark Experiments
We also re-examine all methods on existing benchmark datasets.